OpenStudy (anonymous):

how to show 1/n is "little oh" of 1/(ln(n))? i.e 1/n=o(1/(ln(n)))

OpenStudy (anonymous):

You need to show that \[ \lim_{x\to \infty } \, \frac {\frac 1 x} {\frac 1 {\ln x}}=0 \]

OpenStudy (anonymous):

Which is \[ \lim_{x\to \infty } \, \frac {\ln x}{x}=0 \] Use L'hospital's rule to prove the above limit.